OpenStudy (anonymous):

I tried solving this problem but it kept equaling 0. Can someone help please??? (√6)^(8x)=216(x-3)

OpenStudy (lgbasallote):

\(\sqrt 6 = 6^{1/2}\) \[216 = 6^3\] does that help?

OpenStudy (anonymous):

Hi IGBII

OpenStudy (anonymous):

what does x equal?

OpenStudy (lgbasallote):

remember those previous problems about equaling bases?

OpenStudy (anonymous):

yes, are these the equal bases?

OpenStudy (lgbasallote):

look at my first post..they are both 6 as base right?